OpenStudy (anonymous):

two ships leave port one sailing east and the other south . sometime later they are 17 mi apart, with the eastbound ship 7 mi farther from port than the southbound ship. how far is each from port ?

OpenStudy (mertsj):



OpenStudy (mertsj):

Pythagorean theorem

OpenStudy (mertsj):

\[x^2+(x+7)^2=17\]

OpenStudy (anonymous):

i got : x^2+x^2+49=17 (x+7)(x+7)=17

OpenStudy (anonymous):

where do i go from here @Mertsj

OpenStudy (mertsj):

(x+7)(x+7) is NOT x^2+49 Use FOIL

OpenStudy (anonymous):

?

OpenStudy (mertsj):